OpenStudy (anonymous):

n/3 - 2= 4 How doo i doo thiss?

OpenStudy (anonymous):

(n)/(3)-2=4 Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides. (n)/(3)=2+4 Add 4 to 2 to get 6. (n)/(3)=6 Multiply each term in the equation by 3. (n)/(3)*3=6*3 Simplify the left-hand side of the equation by canceling the common factors. n=6*3 Multiply 6 by 3 to get 18. n=18

OpenStudy (anonymous):

Thank youu!!

OpenStudy (anonymous):

so what if it was like (x)/(3)=7 ? Just time 3 &7?